OpenStudy (anonymous):

PLEASE HELP!!! Solve the equation. sin^2 x + sin x = 2

zepdrix (zepdrix):

Confused on how to start this one? :o

OpenStudy (anonymous):

Yeah...

zepdrix (zepdrix):

Would you know how to solve this equation if you given it? \(\large\rm u^2+u=2\) solve for u?

OpenStudy (anonymous):

Oh! Use it like the quadratic formula?

zepdrix (zepdrix):

Yessss. Good good good.

OpenStudy (anonymous):

So, then I'd subtract two from both sides.. And for my problem I'd treat sin as x

zepdrix (zepdrix):

Well we already have x in the problem, maybe us another letter? :3

zepdrix (zepdrix):

Treat \(\rm\sin x\) as \(\rm u\) maybe? :)

OpenStudy (anonymous):

so u^2 + u = -2

OpenStudy (anonymous):

And it would factor out to (u-1) and (u+2)

zepdrix (zepdrix):

u^2+u-2=0 is what you meant to write, yes? Ok factors look good!

OpenStudy (anonymous):

Yeah, thats what I meant.

OpenStudy (anonymous):

So, then U = -2 and U = 1

OpenStudy (anonymous):

But I don't understand how to use that

zepdrix (zepdrix):

Ok great, let's "undo" the replacement we made earlier.

OpenStudy (anonymous):

sinx = -2 and sinx = 1 ?

zepdrix (zepdrix):

Yessss. Can you think of angles that would give us these values? If they exist.

OpenStudy (anonymous):

I can't think of any..

zepdrix (zepdrix):

sine is your y-coordinate. y-coordinate = 1 at which angle? :)

zepdrix (zepdrix):

the full length

OpenStudy (anonymous):

Oh! sin x = 1 at pi/2

zepdrix (zepdrix):

Ok good :)

OpenStudy (anonymous):

So if sin x = -2, that's just showing that it makes more revolutions?

zepdrix (zepdrix):

Hmm not quite :) Recall that your sine function can only output values between -1 and 1. So -2 is outside of that range. So therefore sinx=-2 gives no solution.

OpenStudy (anonymous):

Wait..

OpenStudy (anonymous):

But sin can't be more than one or less than negative one.. right?

OpenStudy (anonymous):

So would the other answer be extranous?

zepdrix (zepdrix):

ya sorry i spilled the beans :3 bahhh i gotta type slower, sorry

OpenStudy (anonymous):

It's all good! I'm just really slow sometimes. Thanks for all your help!

zepdrix (zepdrix):

extraneous is when we find a solution, and then discover that it doesn't work for the original problem. In this case, we are unable to find a solution. So it's a little different than extraneous, but yes same type of idea ^^

OpenStudy (anonymous):

Oh! So extraneous is when we plug it back in, but it doesn't work?

zepdrix (zepdrix):

yes

OpenStudy (anonymous):

Thank you much!

zepdrix (zepdrix):

oh oh oh

zepdrix (zepdrix):

So yes, pi/2 is a solution. But what happens when we spin another time around the circle, landing up at pi/2 again. This new angle would be pi/2 + 2pi. This also gives us 1 when we take the sine of it, right?

OpenStudy (anonymous):

so it would be 90 and 450 degrees?

zepdrix (zepdrix):

Well, what if we spun around the circle twice and landed at the same spot, 810 degree angle should work also, yes?

OpenStudy (anonymous):

so i would add + 2pi n?

zepdrix (zepdrix):

We have to allow for rotations. (Unless the question specified an interval)

zepdrix (zepdrix):

Yes, good. \(\large\rm \sin(x)=1\) therefore \(\large\rm x=\frac{\pi}{2}+2\pi n\) Where n is an integer.

zepdrix (zepdrix):

Or if you prefer degrees, x=90+360n

OpenStudy (anonymous):

Yay! I finally get it now!!

zepdrix (zepdrix):

yay \c:/